A motorist traveling at 10m/s, was able to bring his car to rest in a distance of 10m. If he had been traveling at 30m/s, in what distance could he bring his cart to rest using the same breaking force?

Grade:12

2 Answers

DINESH HIREGANGE
10 years ago
90m is the answer
DINESH HIREGANGE
10 years ago
v2 = u2 + 2as 0 = 100 + 20a a = -5m/s2 Again, v2 = u2 + 2as 0 = 900 -10s -900/-10 = s s = 90m